### v2

# Performance Rating Equilibrium

Abstract
In this note, I introduce a novel performance rating system called Performance Rating Equilibrium (PRE). A PRE is a vector of hypothetical ratings for each player, such that if these ratings were each player’s initial rating at the start of a tournament, scoring the same points against the same opponents would leave each player’s initial rating unchanged. In other words, all players’ initial ratings perfectly predict their actual scores in the tournament. This property, however, does not hold for the well-known Tournament Performance Rating. PRE is defined as a fixed point of a multidimensional rating function. I show that such a fixed point, and hence a PRE, exists under mild conditions. I provide an implementation of PRE along with several empirical applications. PREs have broad applicability, from sports competitions to the evaluation of large language models. JEL: Z20, D63
Keywords: Tournament performance rating, Elo rating, chess
1 Introduction
Rating and ranking systems are widely studied across various fields, including economics, social choice, sports, and computer science, to assess individual performance or rank objects. One prominent approach, the Elo rating system, originally developed for chess by Elo (1978), has been applied in diverse domains, including association football (FIFA, 2018; Csató, 2021), tennis (Williams et al., 2021), college rankings (Avery et al., 2012), and large language model evaluation (Zheng et al., 2024). [^2] Alternative systems, such as the Glicko rating system (Glickman, 1995), may also be used for ratings in these areas.
In a chess tournament, players’ scores against opponents with specific ratings are used to calculate their Tournament Performance Rating (TPR). The TPR is the hypothetical rating $R$ that would remain unchanged after the tournament. However, because TPR is calculated directly from opponents’ pre-tournament ratings, it does not take into account their actual performances in the event.
Here, I introduce the concept of Performance Rating Equilibrium (PRE), represented by a vector $r^{*}$ of hypothetical ratings. A PRE is defined as a fixed point of a multidimensional rating function: if each player $i$ began the tournament with an initial rating $r^{*}_{i}$, then achieving the same points against the same opponents would leave each player’s initial rating $r^{*}_{i}$ unchanged. In other words, these initial ratings perfectly predict each player’s final score, and I refer to each player’s rating at a PRE as a perfect performance rating (PPR).
Landau (1895, 1914) introduced one of the earliest score-based ranking systems inspired by round-robin chess tournaments. Iterative methods for ranking, whether of players or political candidates, date back to at least Wei (1952); Kendall (1955), who extended an approach by Wei (1952). For some selected further developments, see Keener (1993); Laslier (1997); Herings et al. (2005); Brozos-Vázquez et al. (2008); González-Díaz et al. (2014); Csató (2015, 2017), as well as the influential PageRank algorithm introduced by Brin and Page (1998). Herings et al. (2005) propose a new method for ranking the nodes in directed graphs, while González-Díaz et al. (2014) provide an axiomatic analysis of several ranking methods. [^3] Rubinstein (1980) provides an axiomatization of (non-iterative) score-based ranking. In addition, Csató (2017) explores Swiss-system team tournaments, proposing various score-based ranking methods and applying them to the European Chess Team Championships. [^4] For a broader, growing literature on fairness in sports, see, e.g., Brams and Ismail (2018); Brams et al. (2018); Csató (2021); Lambers and Spieksma (2021); Palacios-Huerta (2023); Devriesere et al. (2024), and the references therein.
The early pioneering work by Kendall (1955) focused on recursive rankings that incorporate both the players’ current performances (measured by scores) and the performances of their opponents in the tournament. To my knowledge, the study by Brozos-Vázquez et al. (2008) is the first to use performance ratings, called “Recursive Performance,” that account for the opponents’ current performance ratings rather than only current scores in their ranking method. There are three main differences between Recursive Performance and the current study. First, Brozos-Vázquez et al.’s main focus is on ranking participants, whereas the focus here is on finding ratings that perfectly predict the actual performances of players. Second, although Recursive Performance offers an improvement over the standard TPR, it, like TPR, does not generally yield a fixed point of the rating function, so these concepts are not in equilibrium as defined here. [^5] The equilibrium Lastly, Recursive Performance ratings are calculated using average opponent ratings with a corrective factor based on score. Although this approach approximates actual performance ratings reasonably well when rating gaps are small, it is well known that it can lead to paradoxical outcomes, such as a player’s performance rating decreasing after winning against a low-rated opponent.
2 Setup
Let $T=(N,r,k,m,f)$ denote a tournament, where $N=\{1,2,\dots,n\}$ is the set of players. The non-negative vectors $r\in\mathbb{R}^{n}_{+}$, $k\in\mathbb{N}^{n}_{+}$, and $m\in\mathbb{R}^{n}_{+}$ denote the initial ratings, the number of games played by the players, and their final scores, respectively, with $0\leq m_{i}\leq k_{i}$ for each $i\in N$. Here, $m_{i}$ denotes $i$’s (final) score in $k_{i}$ games. The function $f:\mathbb{R}^{n}_{+}\rightarrow\mathbb{R}^{n}_{+}$ is called a rating function. A vector of ratings $x^{*}\in\mathbb{R}^{n}_{+}$ is called a Performance Rating Equilibrium (PRE) if $f(x^{*})=x^{*}$. [^6] Here, the term ‘equilibrium’ is used in the sense of a steady state; it is an interactive but a non-choice-based concept.
Let $[f(\cdot)]_{a}^{b}=\min(\max(f(\cdot),a),b)$ denote a function bounded within the interval $[a,b]$. Each player $i$ plays $k_{i}$ games. Denote $i$’s opponents in these games by $i_{1},i_{2},\dots,i_{k_{i}}$. For some $j$ and $j^{\prime}$, it may be that $i_{j}=i_{j^{\prime}}$, meaning that a player may play more than once against the same opponent. Let
$c=\max_{i\in N}\sum_{j=1}^{k_{i}}r_{i_{j}}$
denote the highest possible value for the sum of ratings of a player’s opponents. The role of $c$ is to provide a bound for the rating function. While the main result holds for any upper bound $c>0$ and any chosen lower bound, the definition of $c$ is arguably a reasonable bound for a rating function based on initial ratings $r$ because, even if a player wins all games, the player’s new rating should not strictly exceed the sum of the ratings of the player’s opponents.
Fix a tournament $T$. Note that Tournament Performance Rating (TPR) is undefined for zero or perfect scores, as Elo himself noted (Elo, 1978, p. 12). Ismail (2023) introduces Complete Performance Rating (CPR), which is defined for all scores, including when $m_{i}=0$ or $m_{i}=k_{i}$.
For $x\in\mathbb{R}^{n}_{+}$ and for every player $i$ with $0<m_{i}<k_{i}$, let $y_{i}$ be the value that solves this equation
$m_{i}=\sum_{j=1}^{k_{i}}\frac{1}{1+10^{\frac{x_{i_{j}}-y_{i}}{400}}}.$ (1)
Define $\text{TPR}_{i}(x)=[y_{i}]_{0}^{c}$. For $m_{i}=0$ or $m_{i}=k_{i}$, define $\text{TPR}_{i}(x)=\left[\text{CPR}_{i}(x)\right]_{0}^{c}$.
The following theorem establishes the existence of PREs in tournaments.
Main Theorem: Let $T=(N,r,k,m,\text{TPR})$ be a tournament. Then, there exists a PRE in $T$.
Proof.
I show that the function TPR has a fixed point $x^{*}\in\mathbb{R}^{n}_{+}$ such that $\text{TPR}(x^{*})=x^{*}$. Define the set $C=\{x\in\mathbb{R}_{+}^{n}\mid x_{i}\leq c,\,i\in N\}$. $C$ is compact, as it is closed and bounded.
First, I show that by the Intermediate Value Theorem, for each $x\in C$ and every player $i$ with $0<m_{i}<k_{i}$, there exists a unique $y_{i}$ that satisfies $m_{i}=g_{i}(y_{i})$, where
$g_{i}(y_{i})=\sum_{j=1}^{k_{i}}\frac{1}{1+10^{\frac{x_{i_{j}}-y_{i}}{400}}}.$
Consider the derivative of each term in the summation with respect to $y_{i}$.
$\frac{dg_{i}}{dy_{i}}=\frac{\ln(10)}{400}\frac{10^{\frac{x_{i_{j}}-y_{i}}{400}}}{\left(1+10^{\frac{x_{i_{j}}-y_{i}}{400}}\right)^{2}}.$ (2)
Since this derivative is strictly positive for each term in the summation, the function $g_{i}(y_{i})$ is strictly increasing in $y_{i}$. As $y_{i}$ approaches $\infty$- ∞, $g_{i}(y_{i})$ tends to 0 because $10^{\frac{x_{i_{j}}-y_{i}}{400}}$ tends to $\infty$. Conversely, as $y_{i}$ approaches +$\infty$+ ∞, $g_{i}(y_{i})$ tends to $k_{i}$ since each term $1/(1+10^{\frac{x_{i_{j}}-y_{i}}{400}})$ approaches 1. Since $g_{i}(y_{i})$ is continuous, strictly increasing, and ranges from 00 to $k_{i}$, by the Intermediate Value Theorem, for any $m_{i}$, there exists a unique $y_{i}$ such that $g_{i}(y_{i})=m_{i}$.
Second, I show that TPR is continuous by applying the Implicit Function Theorem. Define the function $F(y_{i},x)$ as follows.
$F(y_{i},x)=\sum_{j=1}^{k_{i}}\frac{1}{1+10^{\frac{x_{i_{j}}-y_{i}}{400}}}-m_{i}.$
The partial derivative of $F$ with respect to $y_{i}$ is shown in Equation 2, and it is the negative of the partial derivative of $F$ with respect to $x_{i_{j}}$. Since the partial derivatives involve exponential functions, they are continuous with respect to both $y_{i}$ and $x_{i_{j}}$.
Now, calculate $\frac{\partial F}{\partial y_{i}}$:
$\frac{\partial F}{\partial y_{i}}=\sum_{j=1}^{k_{i}}\frac{\ln(10)10^{\frac{x_{i_{j}}-y_{i}}{400}}}{400\left(1+10^{\frac{x_{i_{j}}-y_{i}}{400}}\right)^{2}}.$
Since each term in the summation is strictly positive, $\frac{\partial F}{\partial y_{i}}$ is strictly positive.
By the Implicit Function Theorem, there exists an open neighborhood $U$ containing $x$, and a unique continuously differentiable function $y_{i}=\phi_{i}(x)$ defined on $U$, such that $F(\phi_{i}(x),x)=0$ for all $x\in U$.
Third, I extend the function $\phi_{i}(x)$ to $C$. The collection of sets $\{U_{x}\}_{x\in C}$ forms an open cover of $C$. Since $C$ is compact, there exists a finite subcover $\{U_{1},U_{2},\dots,U_{t}\}$ such that
$C\subseteq\bigcup_{j=1}^{t}U_{j}.$
For each $x\in C$, choose any $U_{j}$ containing $x$ and set $\phi_{i}(x)=\phi_{i}^{(j)}(x)$. For any $x\in U_{j}\cap U_{j^{\prime}}$, uniqueness of the solution $y_{i}$ implies that
$\phi_{i}^{(j)}(x)=\phi_{i}^{(j^{\prime})}(x)=y_{i}.$
Thus, the function $\phi_{i}$ is uniquely defined on the domain $C$. Since each $\phi_{i}^{(j)}$ is continuous on $U_{j}$, $\phi_{i}(x)$ is continuous on $C$. As a result, for every player $i$ with $0<m_{i}<k_{i}$, $\text{TPR}_{i}(x)=\phi_{i}(x)$ is continuous.
Since for every player $i$ with $m_{i}=0$ or $m_{i}=k_{i}$, $\text{CPR}_{i}$ is also a continuous function (Ismail, 2023), the function TPR is a continuous function.
Fourth, by definition, for every $x\in C$ and every $i\in N$, $\text{TPR}_{i}(x)\leq c$. Thus, the function TPR maps $C$ into itself. Moreover, $C$ is convex: let $x,x^{\prime}\in C$ and take any $\lambda\in[0,1]$. For each $i$, $0\leq x_{i}\leq c$ and $0\leq x^{\prime}_{i}\leq c$ implies that $0\leq\lambda x_{i}+(1-\lambda)x^{\prime}_{i}\leq c$. Thus, $(\lambda x+(1-\lambda)x^{\prime})\in C$.
Finally, since the function TPR, when restricted to $C$, is continuous and maps $C$ into itself, by Brouwer’s fixed point theorem, there exists $x^{*}\in C$ such that $\text{TPR}(x^{*})=x^{*}$. ∎
This result extends to other continuous rating functions defined on a convex and compact domain, including the Complete Performance Rating (CPR) and simplified versions of TPR that use average ratings (see, e.g., Ismail, 2023).
3 Applications
In a tournament $T$, there may be multiple fixed points or PREs. However, a PRE can be selected in various ways. In the example below, I numerically calculate the fixed point that converges from an initial vector in which all players have the tournament average rating. While equilibrium selection can also be done using other methods, such as starting with the initial ratings vector instead of the average ratings vector, a PRE obtained with the latter approach has the advantage of not depending on the specific ratings that players had at the start of the tournament. In other words, this approach is entirely performance-based, given the tournament average rating. A numerical approximation for calculating PREs is demonstrated in the following implementation: https://github.com/drmehmetismail/Estimated-Performance-Rating. [^7] FIDE (International Chess Federation) uses certain approximations and specific rules when updating ratings and calculating performance ratings. For a more detailed discussion, see Ismail (2023), who gives a definition of performance ratings for all possible scores.
Palma de Mallorca Sharjah
Name Rtg Pts TPR PPR Name Rtg Pts TPR PPR
Aronian 2801 5.5 2821 2857 Vachier-Lagrave 2796 5.5 2823 2860
Jakovenko 2721 5.5 2824 2840 Grischuk 2742 5.5 2828 2852
Nakamura 2780 5 2788 2830 Mamedyarov 2766 5.5 2813 2851
Svidler 2763 5 2779 2815 Nepomniachtchi 2749 5.0 2764 2795
Tomashevsky 2702 5 2788 2813 Nakamura 2785 5.0 2776 2812
Harikrishna 2738 5 2764 2789 Jakovenko 2709 5.0 2781 2818
Ding Liren 2774 5 2768 2783 Adams 2751 5.0 2776 2795
Rapport 2692 5 2758 2743 Ding 2760 5.0 2748 2753
Radjabov 2741 5 2760 2743 Eljanov 2759 4.5 2693 2675
Vachier-Lagrave 2796 4.5 2741 2768 Rapport 2692 4.5 2726 2694
Eljanov 2707 4.5 2724 2706 Li Chao 2720 4.5 2722 2700
Inarkiev 2683 4.5 2735 2699 Vallejo Pons 2709 4.5 2714 2688
Giri 2762 4 2696 2695 Hou Yifan 2651 4.0 2685 2689
Vallejo Pons 2705 4 2682 2643 Aronian 2785 4.0 2696 2693
Li Chao 2741 4 2660 2623 Salem 2656 3.5 2624 2592
Riazantsev 2651 3.5 2641 2622 Tomashevsky 2711 3.5 2629 2591
Hammer 2629 3 2590 2562 Hammer 2628 3.5 2648 2617
Gelfand 2719 3 2582 2555 Riazantsev 2671 3.0 2590 2560
Table 1: Perfect Performance Ratings in the 2017 FIDE Grand Prix Series in Palma de Mallorca and Sharjah 9-round Swiss tournaments
[…]
Note: Rtg = Rating; Pts = Points; TPR = Tournament Performance Rating; PPR = Perfect Performance Rating.
Table 1 presents the PPRs from two top Swiss tournaments: FIDE Grand Prix Series in Palma de Mallorca and Sharjah. In this table, each player’s PPR is the rating $r_{i}$ that satisfies the equation
$m_{i}=\sum_{\begin{subarray}{c}j=1\end{subarray}}^{9}\frac{1}{1+10^{\frac{r_{i_{j}}-r_{i}}{400}}}$
where $m_{i}$ is the player’s final score in the tournament. This equation ensures that players’ final scores are equal to their total expected scores. If players had entered the tournament with their calculated PPRs as initial ratings, their final scores would be precisely the scores predicted by these ratings. In other words, each player’s initial PPR perfectly predicts the player’s final score.
In the Palma de Mallorca Grand Prix, although Jakovenko ranked first by the TPR tiebreak, Aronian’s PPR is higher due to the strong performances of his opponents in this event. For similar reasons, Nakamura’s and Svidler’s PPRs are also significantly higher than their TPRs. In the Sharjah Grand Prix, the differences between PPR and TPR can also be observed for Maxime Vachier-Lagrave and Grischuk.
Rank Name Points PPR Rank Name Points PPR
1 Fischer 18.5 2805 13 Hort 11.5 2556
2 Larsen 15 2669 14 Ivkov 10.5 2525
3 Geller 15 2669 15 Suttles 10 2509
4 Huebner 15 2669 16 Minic 10 2509
5 Taimanov 14 2636 17 Reshevsky 9.5 2493
6 Uhlmann 14 2636 18 Matulovic 9 2477
7 Portisch 13.5 2620 19 Addison 9 2477
8 Smyslov 13.5 2620 20 Filip 8.5 2460
9 Polugaevsky 13 2604 21 Naranja 8.5 2460
10 Gligoric 13 2604 22 Ujtumen 8.5 2460
11 Panno 12.5 2588 23 Rubinetti 5 2350
12 Mecking 12.5 2588 24 Jimenez 5.5 2372
Table 2: Palma de Mallorca Interzonal Tournament (round-robin), 1970
Table 2 shows the PPRs from the 1970 Palma de Mallorca Interzonal Tournament. The Interzonal events were part of the World Chess Championship cycle and are similar to the current Chess World Cup. Since the first official FIDE Elo list was not published until 1971, players in this tournament did not have official ratings. However, based on the first published ratings for the players, it is possible to estimate the tournament’s average rating at approximately 2557. The resulting PPRs are presented in the table.